Status Report

ISS On-Orbit Status 16 Aug 2002

By SpaceRef Editor
August 16, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.   Crew workday schedule was from 8:00 pm EDT last night to 2:10 pm today.  Sleep period ends tonight at 10:00 pm, with a comparatively leisurely day tomorrow.

EVA-7:  Today’s spacewalk did not go quite as planned but accomplished its main objective.  Valery Korzun’s and Peggy Whitson’s egress from the DC-1 “Pirs” was delayed by 1 h 43 min from 3:40 am EDT to 5:23 am, and as a consequence Russian flight controllers deleted two of the three planned tasks from the EVA-7 timeline, viz., (1) replacement of the Kromka payload, and (2) inspection/swabbing of the thruster effluent residues on the SM surface (both tasks are being rescheduled).  Spacewalk duration, constrained by the Orlan LiOH (lithium hydroxide) CO2 filters, could not be extended.  The planned installation of six orbital debris protection panels on the SM’s conical section circumference was completed without problems (17 more panels are to be installed over the next few years).  The EVA ended with hatch closure at 9:48, after a total duration of 4 h 25 min.  It was the 42nd EVA of ISS assembly and maintenance (total accumulated time: 260 h 23 min), the 17th from ISS airlocks (93 h 8 min), and the 7th from the Russian DC-1 airlock.    [After Valery Korzun and Peggy Whitson (on her first EVA), on the original schedule, had donned their Orlans (at 2:00 am EDT), started DC-1 airlock depressurization, completed the 30-min. O2 prebreathe for denitrogenation (to prevent “bends”), and disconnected umbilicals, going to internal power (at 3:41 am) and Orlan-stored oxygen, Caution & Warning (C&W) alarms in both suits indicated that the crucial flow control valves of the primary O2 tanks in both Orlan backpacks were inadvertently misconfigured.  DC-1 had to be repressed (partially, to 550 mmHg; nominal being 760 mmHg), to allow Korzun to egress his Orlan, configure the O2 bottle valves in both suit backpacks, and then repeat the process of suit ingress, depress, prebreathe, etc.]

After Korzun and Whitson had returned on board, the crew reconfigured the ISS, powered up all systems that were deactivated for the EVA and completed post-EVA Orlan operations.  [This included reconfiguring ventilation and installing air ducts in the SM and DC-1, rerouting C&W alarms from the EVA support panels (POV) in DC-1 and PkhO (SM Transfer Compartment) to the SM C&W panel (PSS), and deactivating the Soyuz ASU toilet systems.  Post-EVA Orlan ops followed, such as removal of consumables and batteries from the backpacks and drying out the water lines.  Dryout of the Orlans themselves is also scheduled for tomorrow.]

During the EVA, FE-2 Sergey Treschev closed the Node zenith MPEV (manual pressure equalization valves) to allow repressurization of the ISS “stack” with nitrogen gas (N2) from the Lab and Joint Airlock PCA NIVs (pressure control assembly/nitrogen introduction valves).  Target pressure: 750 mmHg (pressure rise taking 6.5 minutes per mmHg).

Before and after the EVA, Korzun and Whitson, assisted by Treschev, underwent sessions with the standard Russian MO-8 (Body Mass Measurement) and MO-9 (Urolux Urinalysis) biomedical experiments.

Daily routine maintenance was performed today by Peggy (Lab payload status checkup) and Sergey (inspection of BRPK-2 water condensate separator).

ISS flight attitude, first in LVLH (local vertical/local horizontal, “earth-oriented”), was put on Attitude Hold at 1:15 am for the DC-1 depress, returning to LVLH momentum management by the CMGs after the second depress.  At 1:14pm, the station will maneuver back to sun-oriented XPOP (x-axis perpendicular to orbital plane), to remain there under CMG momentum management until EVA-8.

EVA-8, the second Orlan-EVA, while officially still scheduled for next week (8/23), may be slipped by TsUP, with MMT (Mission Management Team) concurrence, a few days to 8/28 to allow adequate time for replanning of goals and timeline.

Today’s targets for the CEO (crew earth observations) program were Industrialized SE Africa (crew was to shoot obliques to capture any aerosols accumulating over the industrial heartland behind the recent cold front), Karachi, Pakistan (Pakistan’s major city and port [5.2 million].  Nadir view), Tunis, Tunisia (nadir view; then views of Etna volcano smoking [east end of Sicily] and lower end of the Italian boot, right of track), Lower Amazon River Basin (the Amazon river mouth remains surprisingly cloud free. Detailed views of the great southern arm of the estuary. Crew was to look down [nadir] and left of track), Lisbon, Portugal (capital city and environs [2.6 million people] right of track, on the big estuary of the Tangos River), Lima, Peru (Lima city just right of track. Then a remarkably clear day expected over a large area of remote western Amazonia and Venezuela, all the way to the Caribbean coast.  Crew was to try a mapping pass of near nadir views of this poorly photographed region [4 minutes from the Andes Mts to the coast]), and Berlin, Germany (this large city [3.47 million] is very dispersed with forests and lakes and is hard to see).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:10 pm EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32-amp mode), on backup pump.  Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed).  U.S. CDRA CO2 scrubber is on Override.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 734, temperature (deg C) — 24.9, ppO2 (mmHg) — 160.5, ppCO2 (mmHg) — 1.8.
  • SM Transfer Compartment:  Pressure (mmHg) — 731, temperature (deg C) — 18.6.
  • FGB Cabin:  Pressure (mmHg) — 740, temperature (deg C) — 20.0.
  • Node:  Pressure (mmHg) — 730.53, temperature (deg C) — 22.6 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 732.80, temperature (deg C) — 25.2, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 732.70, temperature (deg C) — 24.0; shell heater temp (deg C) — 23.6, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.7
  • PMA-2:  Shell heater temp (deg C) — 17.5

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly) were replaced, but some more work needs to be done).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B in Autotrack, mode (solar-tracking).
  • SM batteries: Battery #1 is off-line; all other batteries (7) are in “Complete Charge” mode.
  • FGB batteries:  Battery #5 is off-line; all other batteries (5) are in “Complete Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode (were in Discharge mode during EVA-7).

Thermal Control Systems:

  • Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is operational; PL-2 MDM is Off.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line.
  • State vector — US GPS (SIGI string 1)
  • Attitude — Russian segment
  • Angular rates — US RGA1 (rate gyro assembly 1)

Communications & Tracking Systems:

  • All Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1), in EVA-7 viewing position, with Keep Alive power on both strings (based on MBS).
  • MBS: Keep Alive power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 7:15 am EDT [= epoch]):

  • Mean altitude — 394.1 km
  • Apogee — 405.97 km
  • Perigee — 382.3 km
  • Period — 92.4 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.001747
  • Orbits per 24-hr. day — 15.58
  • Altitude decrease — 200 m (mean) in last 24 hours
  • Solar Beta Angle — -17.0 deg (magnitude increasing)
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 21343
  • Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: ~180 deg, pitch: -5.7 deg., roll: 0 deg]).

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.